OpenStudy (anonymous):

What is the vertex of this equation y = 3(x + 2)^2?

OpenStudy (anonymous):

3x^2 + 12x + 12

OpenStudy (anonymous):

use x = -b/2a

OpenStudy (anonymous):

I think its 12

OpenStudy (anonymous):

if it's written that way, the vertex should be obvious... (-2, 0). the vertex-form for a parabola is: y-k = (x-h)^2 where (h, k) is your vertex.